OpenStudy (carlj0nes):

If I have a pressure vessel to store 5kg of Hydrogen gas at 15 deg c at 10,000psi how much volume would that require?

OpenStudy (michele_laino):

here we can apply the gas state equation: \[\huge PV = nRT\] where \(P\) is pressure of the gas, \(V\) is the volume occupied by the gas, \(n\) is the number of moles of the gas, \(R\) is the gas constant, and \(T\) is the absolute temperature of the gas. Please make sure that all the quantities, in such formula have the units of measure in the same system of units of measure

OpenStudy (michele_laino):

hint: since we consider the hydrogen gas, which is composed by \(H_2\) molecules, then the molecular weight is \(1.008 \cdot 2 \approx 2\)

OpenStudy (carlj0nes):

@Michele_Laino does this look correct. The ideal gas equation is: pv=nRT Pressure, p 6.8948e+7 pa Volume, V ? Number of moles, n, 5000g/(1.008x2) The gas constant, R, The SI value for R is 8.31441 J K-1 mol-1. The temperature, T =288.15K pv=mass(g)/(mass of 1 mole (g)) x RT 6.8948e+7 v=5000/((1.008x2)) x 8.31441 x 288.15 v=(5000/((1.008x2)) x 8.31441 x 288.15)/(6.8948e+7 ) m^3 =0.8619m^3